OpenStudy (bananas1234):

Solve 2x^2 = 72 {±18} {±36} {±6} {±8}

syed98 (syedmohammed98):

2x^2 = 72 x^2 = 72/2 x^2 = 36 x = ?

OpenStudy (bananas1234):

so is it division all the way?

syed98 (syedmohammed98):

when u take the square root of x u should also take the square root of the answer and .the answer is sqaure root of 36

OpenStudy (bananas1234):

ohhh ok so the square root of 36 is 6 so would the answer be {±6}?

syed98 (syedmohammed98):

yup =)

OpenStudy (bananas1234):

what does the ± symbol mean?

syed98 (syedmohammed98):

um it shows that +6 multiplied by +6 gives you 36 whereas -6 multiplied by -6 also gives you 36 =)

OpenStudy (bananas1234):

oh i see, Thanks!

syed98 (syedmohammed98):

ur welcome